OpenStudy (anonymous):

Help! Determine the value of y for the system of equations. x − 3y = 4 2x + y = 8

OpenStudy (anonymous):

I'd cancel x and get -2x + 6 = -8 2x + y = 8 (Upon multiplying by -2) 7x = 0 x = 7, but that's not one of my answers.

OpenStudy (anonymous):

H-help ;-;

OpenStudy (anonymous):

@GoldPhenoix

OpenStudy (goldphenoix):

Sorry. Cannot help you. i don't know how to do this. =/ Sorry for breaking your heart.

OpenStudy (anonymous):

Don't go breakin' my. Don't go breakin' my! D-don't go breakin' my heart!

hartnn (hartnn):

"I'd cancel x and get <-----correct -2x + 6y = -8 <----NOTE 2x + y = 8 (Upon multiplying by -2) " now ADD those 2 equations... -2x+2x+6y+y = -8+8 you can get y from here easily :)